# ICPO: Intrinsic Confidence-Driven Group Relative Preference Optimization for Efficient Reinforcement Learning

### v3

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) demonstrates significant potential in enhancing the reasoning capabilities of Large Language Models (LLMs). However, existing RLVR methods are often constrained by issues such as coarse-grained rewards, reward noise, and inefficient exploration, which lead to unstable training and entropy collapse. To address this challenge, we propose the Intrinsic Confidence-Driven Group Relative Preference Optimization method (ICPO). The intuition behind it lies in the fact that the probabilities of an LLM generating different responses can inherently and directly reflect its self-assessment of the reasoning process. Inspired by the idea of preference modeling, ICPO calculates a preference advantage score for each response by comparing the relative generation probabilities of multiple responses under the same input prompt, and integrates this score with verifiable rewards to guide the exploration process. We have discovered that the preference advantage score not only alleviates the issues of coarse-grained rewards and reward noise but also effectively curbs overconfident errors, enhances the relative superiority of undervalued high-quality responses, and prevents the model from overfitting to specific strategies, thereby facilitating more thorough exploration. Comprehensive experiments across four general-domain benchmarks and three mathematical benchmarks demonstrate that ICPO steadily boosts reasoning compared to GRPO.
@@ 1 Introduction
Large-scale reinforcement learning with verifiable rewards (RLVR) has emerged as a prevailing paradigm for enhancing the reasoning capabilities of LLMs (Open-Reasoner-Zero-2025; Deepcoder-2025; DeepSeek-R1-2025). Unlike reinforcement learning from human feedback (RLHF), RLVR eliminates reliance on subjective human judgments or complex learned reward models by directly employing rule-based reward functions to provide explicit feedback signals for model optimization, thereby establishing a strong coupling between the reward signal and "the correctness of the final answer." This paradigm has given rise to a series of efficient and scalable RLVR training algorithms, such as GRPO (DeepSeek-Coder-2024) and DAPO (DAPO-2025). Despite these significant advances, such methods still face several challenging issues.
@@
To address these challenges, we propose the Intrinsic Confidence-Driven Group Relative Preference Optimization (ICPO), which leverages the model’s inherent self-assessment capability to compensate for the deficiencies of verifiable external rewards. The key intuition lies in the fact that the probability distribution generated by a LLM when producing different reasoning responses essentially represents an implicit self-assessment of the model’s confidence level in its own reasoning process. A higher generation probability indicates that the model deems the reasoning path as highly correct and possesses absolute confidence in it; however, this often corresponds to the model’s path dependency on familiar patterns, potentially resulting in an inertial output tailored to simple scenarios. Conversely, a lower generation probability may actually stem from the model’s attempts to reason through complex or rare samples, even when the model itself exhibits a lack of confidence. This probabilistic preference constitutes a fine-grained signal that can also reflect the effectiveness of policy optimization. Therefore, we can leverage this intrinsic preference as an auxiliary signal to guide policy learning.
Specifically, ICPO draws on the modeling approach of pairwise preferences in direct preference optimization algorithms. For multiple in-group responses generated in response to the same input prompt, it first sorts them in ascending order based on their generation probabilities. Subsequently, it forms pairwise response sets by combining the responses two by two within the group. Then, by comparing the probabilities of these pairwise responses, it calculates a preference advantage score for each response that reflects its relative superiority or inferiority within the group. During the optimization process, ICPO deeply integrates the preference advantage score with traditional verifiable external rewards. On one hand, preference advantage score serves as a supplementary signal, providing fine-grained guidance on the relative merits of intermediate reasoning actions when verifiable rewards are coarse-grained. On the other hand, its stability filters out assessment noise and biases present in verifiable rewards, thereby amplifying the relative value differences between high-quality and low-quality responses. Moreover, this design effectively addresses two typical failure modes in reasoning-based reinforcement learning: (1) It can precisely identify and suppress overconfident errors in the model—that is, the model generates seemingly plausible yet actually erroneous reasoning paths with high probability, while simultaneously enhancing the relative advantage of undervalued high-quality responses; (2) By continuously providing in-group comparison signals, it sustains the exploratory drive of the policy, thereby avoiding training instability and entropy collapse caused by ambiguous or coarse-grained external rewards.
@@
  • Comprehensive experiments across seven benchmarks demonstrate that ICPO consistently outperforms baseline methods (e.g., GRPO) across diverse foundational model architectures including Qwen, Llama, and Gemma, demonstrating superior cross-model generalizability and task transferability.
@@ 2.1 RLVR
Reinforcement Learning with Verifiable Rewards (RLVR), as a robust alternative to Reinforcement Learning from Human Feedback (RLHF), has been demonstrated to effectively enhance model reasoning capabilities (Ganqu-CuiLifan-Yuan-2025; TÜLU3-2025; Deepscaler-2025). While RLHF (Long-Ouyang-2022; Deborah-Cohen-2022; Leo-Gao-2023) leverages human preference data to train reward models—yielding notable improvements—it incurs substantial resource costs due to heavy reliance on manual annotation (Llama2-2023). In contrast, RLVR employs rule-based verification functions (e.g., exact answer matching (Mimo-2025; OpenAI-o1-2024)) to provide reward signals, circumventing the complexity and potential pitfalls of learned reward models. However, this approach remains limited to domains with precise verification tools (e.g., mathematical reasoning and code generation) and suffers from coarse-grained rewards. Recent studies have explored LLM-based verification (LLM-as-a-Judge) (Dawei-Li-2024; PandaLM-2024; JudgeLM-2025) to extend RLVR to open-ended question-answering scenarios. Yet, due to uncertainty in LLM outputs and hallucination issues, the reward signals exhibit low reliability, potentially misleading policy optimization.
@@ 2.2 Self-Reward Optimization
Self-reward optimization via intrinsic signals has emerged as an effective approach to mitigate reliance on extrinsic rewards derived from manual annotation or specialized verification tools. These methods directly extract optimization guidance from the generative process or output features of policy models Weizhe-Yuan-2024; TTRL-2025; Xuandong-Zhao-2025, inspired by observations that LLMs exhibit lower confidence when handling complex problems—optimizing such confidence effectively enhances reasoning capabilities Farquhar-Sebastian-2024; Semantic-Uncertainty-2023; Katie-Kang-2024; Zhewei-Kang-2025. Early studies, such as SPIN Zixiang-Chen-2024 and self-rewarding language models Weizhe-Yuan-2024, leveraged model-generated feedback for subsequent training iterations, while INTUITOR Xuandong-Zhao-2025 employed self-confidence as an intrinsic confidence-based reward. However, such approaches risk limiting exploration Ganqu-Cui-2025; Andreas-Hochlehnert-2025, as excessive reliance on the model’s own epistemic state may trap the policy in self-reinforcing local optima. PLPR RLPR-2025 introduced a stable probability-to-reward conversion method, ensuring precise and consistent rewards derived from LLM intrinsic probabilities. CDE CDE-2025 employs the Perplexity of generated responses as an auxiliary signal to guide policy optimization, offering a novel research pathway for subsequent studies. However, its reliance on absolute assessment of perplexity may lead to suboptimal performance.
@@ 3.1 RL with Intrinsic Confidence-Driven
The core of RLVR lies in replacing traditional RLHF reward models with verifiable, rule-based reward functions (or general LLM scoring), providing the model with explicit binary feedback. This approach eliminates the need for researchers to construct large-scale human preference datasets. Specifically, RLVR optimizes the policy by maximizing the following objective:
@@
where $q$ denotes the input query, $o$ represents the generated output, $\pi_{\text{ref}}$ is the initial reference policy, and $\beta$ controls the KL divergence coefficient to prevent excessive deviation from $\pi_{\text{ref}}$, $v(q,o)$ constitutes a verifiable reward function. Common RLVR algorithms include REINFORCE Ronald-1992, PPO PPO-2017, and GRPO GRPO-2024. GRPO significantly reduces GPU memory consumption and computational costs in training by not needing a value model. Its efficacy has been demonstrated in models such as DeepSeekMath GRPO-2024 and DeepSeek-R1 DeepSeek-R1-2025, establishing it as the prevailing approach in contemporary reinforcement learning. Specifically, for each problem $q$, GRPO samples a set of outputs $\{o_{1},o_{2},…,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
@@
To address the above challenges, we propose Intrinsic Confidence-Driven Group Relative Preference Optimization method (ICPO), as illustrated in Figure 1. This method integrates external reward signals with the model’s self-evaluation of its outputs or reasoning processes. ICPO offers three key benefits: (1) it provides finer optimization guidance under coarse-grained reward conditions; (2) it mitigates random noise in verifiable rewards and accentuates relative value differences between high- and low-quality responses; (3) it reduces prevalent overconfidence errors in reasoning while enhancing the relative superiority of undervalued high-quality responses. The optimization objective of ICPO is:
@@ 3.2 Preference Advantage Scores
where $L_{i}$ represents the effective length of $o_{i}$ (i.e., the number of non-padding tokens), $\pi_{\theta}(o_{t}\mid o_{<t},s)$ denotes the probability of model generating token $o_{t}$ at position $t$, and $\pi_{\theta}(o_{i}|s)$ denotes the sequence-level probability of response $o_{i}$. We normalize probabilities based on the number of tokens to ensure fair comparison across responses of varying lengths and prevent bias toward shorter responses. This ranking structure, which is solely dominated by generation probabilities, is decoupled from external rewards, thereby avoiding noise interference from such rewards.
@@ 3.4 Multi-stage Weight Adjustment
This design originates from the following insight: In early training, high entropy makes it unclear if low-probability responses are worth learning. As training continues and entropy drops, correct and novel low-probability responses emerge, allowing for a progressive increase in their learning momentum. In later training, when most novel samples are learned, remaining low-probability responses may be noise or valueless signals; so reducing intrinsic signals’ influence stabilizes policy learning. Figure 2 (left) displays various weight adjustment strategies, and a detailed analysis of their effects will be presented in the experimental section.
@@ 4 Experiments
This section empirically validates the effectiveness of ICPO in enhancing reasoning abilities of LLMs and explores its potential applications in both verifiable reward and noisy reward domains.
@@ 4.1 Baselines and Setup
We conduct experiments on Gemma2 Gemma2-2024, Llama3.1  llama3-2024 and Qwen2.5 Qwen2.5-2024 series models for fair comparison with existing methods. Unless otherwise specified, experiments are conducted on Qwen2.5-7B-Base. A comprehensive performance evaluation of ICPO method was carried out using three mathematical reasoning benchmarks, namely MATH-500 MATH-500-2024, Minerva Minerva-2022, and AIME24, as well as four general-domain benchmarks, including MMLU-Pro Mmlu-pro-2024, GPQA GPQA-2023, TheoremQA TheoremQA-2023, and WebInstruct. The baseline models include: (1) Base and Instruct models; (2) PRIME Ganqu-CuiLifan-Yuan-2025; (3) SimpleRL-Zoo SimpleRL-Zoo-2025; (4) Oat-Zero Zichen-Liu-2025; (5) TTRL TTRL-2025; (6) General Reasoner General-Reasoner-2025; (7) VeriFree VeriFree-2025; (8) RLPR RLPR-2025; and (9) INTUITOR Xuandong-Zhao-2025. Detailed descriptions of datasets and experimental setups are elaborated upon in the Appendix.
@@ 4.3 Analysis of ICPO
$\Rightarrow$ $\omega$ Warmup decay 57.6 39.4 53.7 78.5 76.2 56.2 23.3 57.3 51.9
@@ 4.4 Natural Scalability
It is noteworthy that GRPO training on Qwen2.5-7B-Instruct shows sustained performance decline in later stages (as shown in Figure 5), whereas ICPO maintains stable exploration. This relates to reduced in-group reward variance caused by the enhanced performance of instruction-tuned models, further validating the robustness and anti-degradation benefits of intrinsic confidence-driven mechanism in coarse-grained reward scenarios.
@@ A.2 Evaluation data
Log-probability: $P_{1}(-0.0135)$, $P_{2}(-0.0583)$, $P_{3}(-0.0216)$, $P_{4}(-0.0090)$, $P_{5}(-0.0407)$
- $S^{p}_{2}=0.4\cdot(\frac{-0.0407}{-0.0583}+\frac{-0.0216}{-0.0583}+\frac{-0.0135}{-0.0583}+\frac{-0.0090}{-0.0583})-0.4\cdot(-0.0090)=0.5852$
- $S^{p}_{5}=0.4\cdot(\frac{-0.0216}{-0.0407}+\frac{-0.0135}{-0.0407}+\frac{-0.0090}{-0.0407})-0.4\cdot(-0.0090)=0.4372$
- $S^{p}_{3}=0.4\cdot(\frac{-0.0135}{-0.0216}+\frac{-0.0090}{-0.0216})-0.4\cdot(-0.0090)=0.4202$
- $S^{p}_{1}=0.4\cdot(\frac{-0.0090}{-0.0135})-0.4\cdot(-0.0090)=0.2704$
- $S^{p}_{4}=-0.4\cdot(-0.0090)=0.0036$
- $R^{final}_{2}=0.1000+1.0\cdot\operatorname{min}(0.5852,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{5}=1.0000+1.0\cdot\operatorname{min}(0.4372,\frac{1.0000}{2.0})=1.4372$
- $R^{final}_{3}=1.0000+1.0\cdot\operatorname{min}(0.4202,\frac{1.0000}{2.0})=1.4202$
- $R^{final}_{1}=0.1000+1.0\cdot\operatorname{min}(0.2704,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{4}=0.1000+1.0\cdot\operatorname{min}(0.0036,\frac{0.1000}{2.0})=0.1036$
- ICPO: $A_{1}(-0.7917)$, $A_{2}(-0.7917)$, $A_{3}(1.2108)$, $A_{4}(-0.8649)$, $A_{5}(1.2376)$
Log-probability: $P_{1}(-0.0386)$, $P_{2}(-0.0387)$, $P_{3}(-0.0224)$, $P_{4}(-0.0086)$, $P_{5}(-0.2568)$
- $S^{p}_{5}=0.4\cdot(\frac{-0.0387}{-0.2568}+\frac{-0.0386}{-0.2568}+\frac{-0.0224}{-0.2568}+\frac{-0.0086}{-0.2568})-0.4\cdot(-0.0086)=0.1186$
- $S^{p}_{2}=0.4\cdot(\frac{-0.0386}{-0.0387}+\frac{-0.0224}{-0.0387}+\frac{-0.0086}{-0.0387})-0.4\cdot(-0.0086)=0.4777$
- $S^{p}_{1}=0.4\cdot(\frac{-0.0224}{-0.0386}+\frac{-0.0086}{-0.0386})-0.4\cdot(-0.0086)=0.3454$
- $S^{p}_{3}=0.4\cdot(\frac{-0.0086}{-0.0224})-0.4\cdot(-0.0086)=0.1569$
- $S^{p}_{4}=-0.4\cdot(-0.0086)=0.0034$
- $R^{final}_{5}=1.0000+1.0\cdot\operatorname{min}(0.1186,\frac{1.0000}{2.0})=1.1186$
- $R^{final}_{2}=0.1000+1.0\cdot\operatorname{min}(0.4777,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{1}=0.1000+1.0\cdot\operatorname{min}(0.3454,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{3}=1.0000+1.0\cdot\operatorname{min}(0.1569,\frac{1.0000}{2.0})=1.1569$
- $R^{final}_{4}=0.1000+1.0\cdot\operatorname{min}(0.0034,\frac{0.1000}{2.0})=0.1034$
- ICPO: $A_{1}(-0.7842)$, $A_{2}(-0.7842)$, $A_{3}(1.2626)$, $A_{4}(-0.8790)$, $A_{5}(1.1847)$
Log-probability: $P_{1}(-0.0135)$, $P_{2}(-0.0583)$, $P_{3}(-0.0216)$, $P_{4}(-0.0090)$, $P_{5}(-0.0407)$
- $S^{p}_{2}=0.4\cdot(\frac{-0.0407}{-0.0583}+\frac{-0.0216}{-0.0583}+\frac{-0.0135}{-0.0583}+\frac{-0.0090}{-0.0583})-0.4\cdot(-0.0090)=0.5852$
- $S^{p}_{5}=0.4\cdot(\frac{-0.0216}{-0.0407}+\frac{-0.0135}{-0.0407}+\frac{-0.0090}{-0.0407})-0.4\cdot(-0.0090)=0.4372$
- $S^{p}_{3}=0.4\cdot(\frac{-0.0135}{-0.0216}+\frac{-0.0090}{-0.0216})-0.4\cdot(-0.0090)=0.4202$
- $S^{p}_{1}=0.4\cdot(\frac{-0.0090}{-0.0135})-0.4\cdot(-0.0090)=0.2704$
- $S^{p}_{4}=-0.4\cdot(-0.0090)=0.0036$
- $R^{final}_{2}=0.1000+1.0\cdot\operatorname{min}(0.5852,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{5}=0.7000+1.0\cdot\operatorname{min}(0.4372,\frac{0.6000}{2.0})=1.0$
- $R^{final}_{3}=1.0000+1.0\cdot\operatorname{min}(0.4202,\frac{1.0000}{2.0})=1.4202$
- $R^{final}_{1}=0.4000+1.0\cdot\operatorname{min}(0.2704,\frac{0.4000}{2.0})=0.60$
- $R^{final}_{4}=0.4000+1.0\cdot\operatorname{min}(0.0036,\frac{0.4000}{2.0})=0.4036$
- ICPO $w$ $R^{Noisy}$: $A_{1}(-0.2556)$, $A_{2}(-1.2578)$, $A_{3}(1.5712)$, $A_{4}(-0.6930)$, $A_{5}(0.6353)$

### v4

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) demonstrates significant potential in enhancing the reasoning capabilities of Large Language Models (LLMs). However, existing RLVR methods are often constrained by issues such as coarse-grained rewards, reward noise, and inefficient exploration, which lead to unstable training and entropy collapse. To address this challenge, we propose the Intrinsic Confidence-Driven Group Relative Preference Optimization method (ICPO). The intuition behind it lies in the fact that the probabilities of an LLM generating different responses can inherently and directly reflect its self-assessment of the reasoning process. Inspired by the idea of preference modeling, ICPO calculates a preference advantage score for each response by comparing the relative generation probabilities of multiple responses under the same input prompt, and integrates this score with verifiable rewards to guide the exploration process. We have discovered that the preference advantage score not only alleviates the issues of coarse-grained rewards and reward noise but also effectively curbs overconfident errors, enhances the relative superiority of undervalued high-quality responses, and prevents the model from overfitting to specific strategies. Comprehensive experiments across four general-domain benchmarks and three mathematical benchmarks demonstrate that ICPO steadily boosts reasoning compared to GRPO.
@@ 1 Introduction
Large-scale reinforcement learning with verifiable rewards (RLVR) has emerged as a prevailing paradigm for enhancing the reasoning capabilities of LLMs (Open-Reasoner-Zero-2025; Deepcoder-2025; DeepSeek-R1-2025). Unlike reinforcement learning from human feedback (RLHF), RLVR eliminates reliance on subjective human judgments or complex learned reward models by directly employing rule-based reward functions to provide explicit feedback signals for model optimization. This paradigm has given rise to a series of efficient and scalable RLVR training algorithms, such as GRPO (DeepSeek-Coder-2024) and DAPO (DAPO-2025). Despite these significant advances, such methods still face several challenging issues.
@@
To address these challenges, we propose the Intrinsic Confidence-Driven Group Relative Preference Optimization (ICPO), which leverages the model’s inherent self-assessment capability to compensate for the deficiencies of verifiable rewards. The key intuition lies in the fact that the probability distribution generated by LLMs when producing different reasoning responses essentially represents an implicit self-assessment of the model’s confidence in its own reasoning. A higher generation probability means the model deems the reasoning path as highly correct and possesses absolute confidence; however, this often corresponds to the model’s path dependency on familiar patterns, potentially leading to inertial outputs for simple scenarios. Conversely, a lower generation probability may actually come from the model’s attempts to reason through complex or rare samples, even when it lacks confidence. This probabilistic preference constitutes a fine-grained signal that can also reflect the effectiveness of policy optimization. Therefore, we can leverage this intrinsic preference as an auxiliary signal to guide policy learning.
Specifically, ICPO draws on the modeling approach of pairwise preferences in direct preference optimization algorithms. For multiple in-group responses generated in response to the same input prompt, it first sorts them in ascending order based on their generation probabilities. Subsequently, it forms pairwise response sets by combining the responses two by two within the group. Then, by comparing the probabilities of these pairwise responses, it calculates a preference advantage score for each response that reflects its relative superiority or inferiority within the group. During the optimization process, ICPO deeply integrates the preference advantage score with traditional verifiable external rewards. This design effectively addresses two typical failure modes in reasoning-based reinforcement learning: (1) It can precisely identify and suppress overconfident errors in the model—that is, the model generates seemingly plausible yet actually erroneous reasoning paths with high probability, while simultaneously enhancing the relative advantage of undervalued high-quality responses; (2) By continuously providing fine-grained comparative signals regarding relative merits within groups, it sustains the exploratory drive of the policy, thereby avoiding training instability and entropy collapse caused by ambiguous or coarse-grained external rewards.
@@
  • Comprehensive experiments on seven benchmarks demonstrate that ICPO consistently outperforms baselines (e.g., GRPO) across various model architectures like Qwen, Llama, and Gemma, demonstrating superior effectiveness and cross-model generalizability.
@@ 2.1 RLVR
Reinforcement Learning with Verifiable Rewards (RLVR), as a robust alternative to Reinforcement Learning from Human Feedback (RLHF), has been demonstrated to effectively enhance model reasoning capabilities (Ganqu-CuiLifan-Yuan-2025; TÜLU3-2025; Deepscaler-2025). While RLHF (Long-Ouyang-2022; Deborah-Cohen-2022; Leo-Gao-2023) leverages human preference data to train reward models—yielding notable improvements—it incurs substantial resource costs due to heavy reliance on manual annotation (Llama2-2023). In contrast, RLVR employs rule-based verification functions (e.g., exact answer matching (Mimo-2025; OpenAI-o1-2024)) to provide reward signals, circumventing the complexity and potential pitfalls of learned reward models. However, this approach remains limited to domains with precise verifiers and suffers from coarse-grained rewards. Recent studies have explored LLM-based verifiers (LLM-as-a-Judge) (Dawei-Li-2024; PandaLM-2024; JudgeLM-2025) to extend RLVR to open-ended question-answering scenarios. Yet, due to uncertainty in LLM outputs and hallucination issues, the reward signals exhibit low reliability, potentially misleading policy optimization.
@@ 2.2 Self-Reward Optimization
Self-reward optimization via intrinsic signals has emerged as an effective approach to mitigate reliance on extrinsic rewards derived from manual annotation or specialized verification tools. These methods directly extract optimization guidance from the generative process or output features of policy models Weizhe-Yuan-2024; TTRL-2025; Xuandong-Zhao-2025, inspired by observations that LLMs exhibit lower confidence when handling complex problems—optimizing such confidence effectively enhances reasoning capabilities Farquhar-Sebastian-2024; Semantic-Uncertainty-2023; Katie-Kang-2024; Zhewei-Kang-2025. Early studies, such as SPIN Zixiang-Chen-2024 and self-rewarding LLMs Weizhe-Yuan-2024, leveraged model-generated feedback for subsequent training iterations, while INTUITOR Xuandong-Zhao-2025 employed self-confidence as an intrinsic confidence-based reward. However, such methods risk limiting exploration Ganqu-Cui-2025; Andreas-Hochlehnert-2025, as excessive reliance on model’s own epistemic state may trap the policy in self-reinforcing local optima. PLPR RLPR-2025 introduced a stable probability-to-reward conversion method. CDE CDE-2025 employs the Perplexity of generated responses as an auxiliary signal to guide policy optimization. However, its reliance on absolute assessment of perplexity may lead to suboptimal performance.
@@ 3.1 RL with Intrinsic Confidence-Driven
The core of RLVR lies in replacing traditional RLHF reward models with verifiable, rule-based reward functions (or general LLM scoring). This approach eliminates the need for researchers to construct large-scale human preference datasets. Specifically, RLVR optimizes the policy by maximizing the following objective:
@@
where $q$ denotes the input query, $o$ represents the generated output, $\pi_{\text{ref}}$ is the initial reference policy, and $\beta$ controls the KL divergence coefficient to prevent excessive deviation from $\pi_{\text{ref}}$, $v(q,o)$ constitutes a verifiable reward function. Common RLVR algorithms include REINFORCE Ronald-1992, PPO PPO-2017, and GRPO GRPO-2024. GRPO significantly reduces GPU memory consumption and training computational costs by not needing the value model. Its efficacy has been validated in DeepSeekMath GRPO-2024 and DeepSeek-R1 DeepSeek-R1-2025, establishing its prominent position in reinforcement learning. Specifically, for each problem $q$, GRPO samples a set of outputs $\{o_{1},o_{2},…,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
@@
To address the above challenges, we propose Intrinsic Confidence-Driven Group Relative Preference Optimization (ICPO), as illustrated in Figure 1. This method integrates external rewards with the model’s self-evaluation of its reasoning processes. ICPO offers three key benefits: (1) it provides finer optimization guidance under coarse-grained reward conditions; (2) it mitigates random noise in verifiable rewards and accentuates relative value differences between high- and low-quality responses; (3) it reduces prevalent overconfidence errors in reasoning while enhancing the relative superiority of undervalued high-quality responses. The optimization objective of ICPO is:
@@ 3.2 Preference Advantage Scores
where $L_{i}$ represents the effective length of $o_{i}$ (i.e., the number of non-padding tokens), $\pi_{\theta}(o_{t}\mid o_{<t},s)$ denotes the probability of model generating token $o_{t}$ at position $t$, and $\pi_{\theta}(o_{i}|s)$ denotes the sequence-level probability of response $o_{i}$. We normalize probabilities based on effective length to ensure fair comparison across responses and prevent bias toward shorter responses. This ranking structure, which is solely dominated by generation probabilities, is decoupled from external rewards, thereby avoiding noise interference from such rewards.
@@ 3.4 Multi-stage Weight Adjustment
This design originates from the following insight: In early training, high entropy makes it challenging to determine whether low-probability responses are worth learning. As training continues and entropy drops, correct and novel low-probability responses emerge, at which point we can incrementally enhance the learning motivation for these responses. In later training, when most novel samples are learned, remaining low-probability responses may be noise or valueless signals; so reducing intrinsic signals’ influence stabilizes policy learning. Figure 2 (left) displays various weight adjustment strategies, with a detailed analysis of their effects in the experimental section. The warmup turning point is set at 0.4 because, in GRPO training, the 20%-40% phase shows suitable entropy reduction with active exploration, making it ideal for introducing strong intrinsic preference rewards. An ablation study on this setting is included in the Appendix.
@@ 4 Experiments
This section empirically validates the effectiveness of ICPO and explores its potential applications in both verifiable reward and noisy reward domains.
@@ 4.1 Baselines and Setup
We conduct experiments on Gemma2 Gemma2-2024, Llama3.1  llama3-2024 and Qwen2.5 Qwen2.5-2024 series models for fair comparison with existing methods. Unless otherwise specified, experiments are conducted on Qwen2.5-7B-Base. A comprehensive performance evaluation of ICPO method was carried out using three mathematical reasoning benchmarks, namely MATH-500 MATH-500-2024, Minerva Minerva-2022, and AIME24, as well as four general-domain benchmarks, including MMLU-Pro Mmlu-pro-2024, GPQA GPQA-2023, TheoremQA TheoremQA-2023, and WebInstruct. The baseline models include: (1) Base and Instruct models; (2) PRIME Ganqu-CuiLifan-Yuan-2025; (3) SimpleRL-Zoo SimpleRL-Zoo-2025; (4) Oat-Zero Zichen-Liu-2025; (5) TTRL TTRL-2025; (6) General Reasoner General-Reasoner-2025; (7) RLPR RLPR-2025; (8) INTUITOR Xuandong-Zhao-2025; and (9) CDE CDE-2025. Detailed descriptions of datasets and experimental setups are elaborated upon in the Appendix.
@@ 4.3 Analysis of ICPO
$\Rightarrow$ $\omega$ Warmup decay 57.6 39.4 54.0 77.8 76.2 56.2 23.3 57.2 51.9
@@ 4.4 Natural Scalability
Notably, GRPO training on Qwen2.5-7B-Inst. shows sustained performance decline in later stages (as shown in Figure 5), whereas ICPO maintains stable exploration. We attribute this to the substantial performance improvement of the instruction-tuned model, which results in smaller relative advantages among different responses within groups. This further validating the robustness and anti-degradation benefits of intrinsic confidence-driven mechanism in coarse-grained reward scenarios.